Using the method of undetermined coefficients, write the trial solution of the equation
d^2y/dx^2 + 2dy/dx +5y= xe^-x cos2x

and hence solve it.
1
Expert's answer
2020-08-19T17:01:34-0400

"\frac{d^2y}{dx^2}+2\frac{dy}{dx}+5y=xe^{-x} \cos2x"

Solution of equation can be found in the form

"y=y_1+y_2"

Solution "y_1" can be found from equation

"y''+2y'+5y=0"

"k^2+2k+5=0\\
D=4-20=-16=16i^2\\
k_1=\frac{-2-4i}{2}=-1-2i\\
k_2=\frac{-2+4i}{2}=-1+2i\\
y_1=e^{-x}(C_1\cos2x+C_2\sin2x)"

Solution "y_2" can be found in form

"y_2=xe^{-x}((ax+b)\cos2x+(cx+m)\sin2x)=\\
=e^{-x}((ax^2+bx)\cos2x+(cx^2+mx)\sin2x))\\
y_2' =e^{-x}\cos2x(x^2(-a+2c)+\\+x(2a-b+2m)+b)+\\
+e^{-x}\sin2x(x^2(-2a-c)+\\+x(-2b+2c-m)+m)\\
y_2'' =e^{-x}\cos2x(x^2(-3a-4c)+\\+x(-4a-3b-4m+8c)-2b+2a+4m)+\\
+e^{-x}\sin2x(x^2(4a-3c)+\\+x(-8a+4b-3m-4c)-4b-2m+2c)\\"

Plug "y_2, y_2', y_2''" into the equation, then

"e^{-x}\cos2x(x^2(-3a-4c)+\\+x(-4a-3b-4m+8c)-2b+2a+4m)+\\
+e^{-x}\sin2x(x^2(4a-3c)+\\+x(-8a+4b-3m-4c)-4b-2m+2c)+\\
+2e^{-x}\cos2x(x^2(-a+2c)+\\+x(2a-b+2m)+b)+\\
+2e^{-x}\sin2x(x^2(-2a-c)+\\+x(-2b+2c-m)+m)+\\
+5e^{-x}((ax^2+bx)\cos2x+(cx^2+mx)\sin2x))=\\
=xe^{-x}\cos2x\\
\cos2x:\\ x^2:-3a-4c-2a+4c+5a=0\\\implies0=0\\
x:-4a-3b-4m+8c+4a-2b+4m+5b=1\\
\implies8c=1, c=\frac{1}{8}\\
x^0:-2b+2a+4m+2b=0\\
\implies2a+4m=0\\
\sin2x:\\
x^2:4a-3c-4a-2c+5c=0\\
\implies0=0\\
x:-8a+4b-3m-4c-4b+\\+4c-2m+5m=0\\
\implies a=0\\
x^0:-4b-2m+2c+2m=0\\
\implies-4b+2c=0\\
b=\frac{1}{16}\\
m=0"

Then

"y_2=e^{-x}(\frac{1}{16}x\cos2x+\frac{1}{8}x^2\sin2x)"

The solution of the equation is

"y=e^{-x}(C_1\cos2x+C_2\sin2x)+\\
+e^{-x}(\frac{1}{16}x\cos2x+\frac{1}{8}x^2\sin2x)"